Brad is 7 years older than Pedro. Six years ago Pedro's age was three-fourths of Brad's age at that time. Find the present ages of Brad and Pedro.
Answer by lwsshak3(11628)

Brad is 7 years older than Pedro. Six years ago Pedro's age was three-fourths of Brad's age at that time. Find the present ages of Brad and Pedro.
**
let x=Pedro's present age
x+7=Brad's present age
6 yrs ago:
x-6=Pedro's age
x+7-6=x+1=Brad's age
x-6=(3/4)(x+1)
LCD:4
4x-24=3x+3
x=27
x+7=34
Pedro's present age=27
Brad's present age=34